Solve the system by the elimination method.
2a + 3b = 6
5a + 2b - 4 = 0
I still don't understand it. Thank you!

Answer by stanbon(75887)

Solve the system by the elimination method.
2a + 3b = 6
5a + 2b - 4 = 0
------
multiply thru the top equation by 2
Multiply thru the bottom equation by 3
--------------------------------------
4a + 6b = 12
15a+ 6b = 12
---------------
Subtract and solve for "a":
11a = 0
a = 0
---------
Solve for "b" using 2a + 3b = 6
2*0 + 3b = 6
b = 2
--------------
Cheers,
Stan H.
==============